OpenStudy (anonymous):

Solve The Following Equation 2(h-8)-h=h-16 A. h=8 B. h=-8 C. infinitely many solutions D. no solution

OpenStudy (anonymous):

\[2h-16-h=h-16\]

OpenStudy (anonymous):

\[h-16=h-16\]

OpenStudy (anonymous):

No solution

OpenStudy (anonymous):

D